A girl's age is half of her mother's age,her mother is 5 years younger than her father,the total of their ages =100
What is the age of the girl?
What is the age of the girl's mother?
What is the age of the girl's father?

Answer by Alan3354(69443)

A girl's age is half of her mother's age,her mother is 5 years younger than her father,the total of their ages =100
==============
F = M + 5
G = M/2
M = M
----------------- Add
M+5 + M/2 + M = 100
5M/2 = 95
M = 38
etc